OpenStudy (anonymous):

I need help understanding if I have started on the right path to solving this problem.

OpenStudy (phi):

why did you use K=2 ? wouldn't it be K= 1/5, so that the prob sum to 1?

OpenStudy (phi):

so your table should have 1/5 everywhere you have a 2, and 2/5 where you have a 4 the marginal distributions are the sum of the rows (for x) and columns (for y)

OpenStudy (phi):

next you need to calculate the mean for x and for y

OpenStudy (phi):

Aren't there formulas for figuring out the mean of U using the mean of X and mean of Y ? ditto for the variance

OpenStudy (phi):

I would find the mean of X using your table then find the mean of Y mean U = mean(X)+ mean(Y) var(U)= var(X)+var(Y)

OpenStudy (phi):

yes, the wording is not clear to me, but I think they want the distributions for X and Y then they want you to use that info to find the mean and variance of the variables U and W

OpenStudy (phi):

for example mean(X)= 0 * 3/5 + 1 * 2/5 = 2/5